# Human Cognition Inspired RAG with Knowledge Graph for Complex Problem Solving

### v1

@@ Abstract
Large language models (LLMs) have demonstrated transformative potential across various domains, yet they face significant challenges in knowledge integration and complex problem reasoning, often leading to hallucinations and unreliable outputs. Retrieval-Augmented Generation (RAG) has emerged as a promising solution to enhance LLMs accuracy by incorporating external knowledge. However, traditional RAG systems struggle with processing complex relational information and multi-step reasoning, limiting their effectiveness in advanced problem-solving tasks. To address these limitations, we propose CogGRAG, a cognition inspired graph-based RAG framework, designed to improve LLMs performance in Knowledge Graph Question Answering (KGQA). Inspired by the human cognitive process of decomposing complex problems and performing self-verification, our framework introduces a three-stage methodology: decomposition, retrieval, and reasoning with self-verification. By integrating these components, CogGRAG enhances the accuracy of LLMs in complex problem solving. We conduct systematic experiments with three LLM backbones on four benchmark datasets, where CogGRAG outperforms the baselines.
Human Cognition Inspired RAG with Knowledge Graph for Complex Problem Solving
Yao Cheng, Yibo Zhao, Jiapeng Zhu, Yao Liu, Xing Sun, Xiang Li
1 Introduction
@@
As a foundational technology for artificial general intelligence (AGI), large language models (LLMs) have achieved remarkable success in practical applications, demonstrating transformative potential across a wide range of domains Touvron et al. (2023); AI@Meta (2024); Yang et al. (2024). Their ability to process, generate, and reason with natural language has enabled significant advancements in areas such as machine translation Zhu et al. (2023), text summarization Basyal and Sanghvi (2023), and question answering Pan et al. (2023). Despite their impressive performance, LLMs still face significant limitations in knowledge integration beyond their pre-trained data boundaries. These limitations often lead to the generation of plausible but factually incorrect responses, a phenomenon commonly referred to as hallucinations, which undermines the reliability of LLMs in critical applications.
@@
To mitigate hallucinations, Retrieval-Augmented Generation (RAG) Niu et al. (2023); Gao et al. (2023); Edge et al. (2024) has emerged as a promising paradigm, significantly improving the accuracy and reliability of LLM-generated contents through the integration of external knowledge. However, while RAG successfully mitigates certain aspects of hallucination, it still exhibits inherent limitations in processing complex relational information. As shown in Figure 1 (a), the core limitation of standard RAG systems lies in their reliance on vector-based similarity matching, which processes knowledge segments as isolated units without capturing their contextual interdependencies or semantic relationships Jin et al. (2024b). Consequently, traditional RAG implementations are inadequate for supporting advanced reasoning capabilities in LLMs, particularly in scenarios requiring complex problem-solving, multi-step inference, or sophisticated knowledge integration.
@@
Recently, graph-based RAG Edge et al. (2024); Ma et al. (2024); Jin et al. (2024b); Mavromatis and Karypis (2024) has been proposed to address the limitations of conventional RAG systems by incorporating deep structural information from external knowledge sources. These approaches typically utilize knowledge graphs (KGs) to model complex relation patterns within external knowledge bases, employing structured triple representations (<entity, relation, entity>) to integrate fragmented information across multiple document segments. While graph-based RAG has shown promising results in mitigating hallucination and improving factual accuracy, several challenges remain unresolved:
@@
$\bullet$ Complex Problem Reasoning. Complex problems cannot be resolved through simple queries; they typically require multi-step reasoning to derive the final answer. While previous work has sought to enhance reasoning capabilities through methods such as chain-of-thought (CoT) reasoning or multi-hop information retrieval, these search-based and iterative approaches still face inherent disadvantages Jin et al. (2024b); Mavromatis and Karypis (2024). As shown in Figure 1 (b), each step in the iterative process relies on the result of the previous step, indicating that errors occurred at previous steps can propagate to subsequent steps.
@@
$\bullet$ Hallucinations. Despite the integration of external knowledge sources, large language models (LLMs) remain prone to generating inaccurate or fabricated responses when confronted with retrieval errors or insufficient knowledge coverage. However, prior research has lacked mechanisms for self-verification, which undermines the reliability of graph-based RAG in real-world applications.
To address these challenges, we propose CogGRAG, a Cognition inspired Graph RAG framework, designed to enhance the complex problem-solving capabilities of LLMs in Knowledge Graph Question Answering (KGQA) tasks. The framework introduces a three-stage methodology: (1) Decomposition, inspired by human problem-solving strategies, breaks down complex problems top-down into smaller, simpler sub-problems and forms a mind map. (2) Retrieval, built on all the decomposed sub-problems, this stage extracts multi-level structured information from external knowledge sources using knowledge graphs. This stage operates at both local and global levels: local retrieval identifies relevant information for individual sub-problems, while global retrieval establishes connections between multiple sub-problems, ensuring detailed and comprehensive knowledge integration. (3) Reasoning with Self-Verification, mimicking human self-reflection, this stage evaluates the reasoning process and verifies the accuracy of intermediate and final results. CogGRAG emulates human cognitive processes in tackling complex problems by employing a top-down decomposition strategy, comprehensive knowledge retrieval, and bottom-up reasoning with validation. By integrating these components, CogGRAG enhances the accuracy and reliability of LLM-generated content on KGQA tasks.
2 Related Work
@@
Figure 2: The overall process of CogGRAG. Given a target question $Q$, CogGRAG first prompts the LLM to decompose it into a hierarchy of sub-problems in a top-down manner, constructing a structured mind map. Subsequently, CogGRAG prompts the LLM to extract both local and global key information from these sub-problems. Finally, CogGRAG guides the LLM to perform bottom-up reasoning and verification based on the mind map and the retrieved knowledge, until the final answer is derived.
@@
Reasoning with LLM Prompting. Recent advancements in prompt engineering have demonstrated that state-of-the-art prompting techniques can significantly enhance the reasoning capabilities of LLMs on complex problems Wei et al. (2022); Yao et al. (2024); Besta et al. (2024). Chain of Thought (CoT Wei et al. (2022) explores how generating a chain of thought—a series of intermediate reasoning steps—significantly improves the ability of large language models to perform complex reasoning. Tree of Thoughts (ToT) Yao et al. (2024) introduces a new framework for language model inference that generalizes the popular Chain of Thought approach to prompting language models, enabling exploration of coherent units of text (thoughts) as intermediate steps toward problem-solving. The Graph of Thoughts (GoT) Besta et al. (2024) models the information generated by a LLM as an arbitrary graph, enabling LLM reasoning to more closely resemble human thinking or brain mechanisms. However, these methods remain constrained by the limitations of the model’s pre-trained knowledge base and are unable to address hallucination issues stemming from the lack of access to external, up-to-date information.
@@
Knowledge Graph Augmented LLM. KGs Vrandečić and Krötzsch (2014) offer distinct advantages in enhancing LLMs with structured external knowledge. Early graph-based RAG methods Edge et al. (2024); He et al. (2024); Hu et al. (2024); Wu et al. (2023); Jin et al. (2024a) demonstrated this potential by retrieving and integrating structured, relevant information from external sources, enabling LLMs to achieve superior performance on knowledge-intensive tasks. However, these methods exhibit notable limitations when applied to complex problems. Recent advancements have introduced methods like chain-of-thought prompting to enhance LLMs’ reasoning on complex problems Sun et al. (2023); Ma et al. (2024); Jin et al. (2024b). Think-on-Graph Sun et al. (2023) introduces a new approach that enables the LLM to iteratively execute beam search on the KGs, discover the most promising reasoning paths, and return the most likely reasoning results. ToG-2 Ma et al. (2024) achieves deep and faithful reasoning in LLMs through an iterative knowledge retrieval process that involves collaboration between contexts and the KGs. GRAPH-COT Jin et al. (2024b) propose a simple and effective framework to augment LLMs with graphs by encouraging LLMs to reason on the graph iteratively. However, iterative reasoning processes are prone to error propagation, as errors cannot be corrected once introduced.
3 Preliminaries
3.1 Problem Definition
@@
In this paper, we focus on the KGQA task. The input of the task is the question $Q$, and the output is the answer $A$. Given the input question $Q$, we aim to design a graph-based RAG framework with an LLM backbone $p_{\theta}$ to enhance the complex problem-solving capabilities and generate the answer $A$.
3.2 Mind Map

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential across various domains. However, they often struggle with integrating external knowledge and performing complex reasoning, leading to hallucinations and unreliable outputs. Retrieval Augmented Generation (RAG) has emerged as a promising paradigm to mitigate these issues by incorporating external knowledge. Yet, conventional RAG approaches, especially those based on vector similarity, fail to effectively capture relational dependencies and support multi-step reasoning. In this work, we propose CogGRAG, a human cognition-inspired, graph-based RAG framework designed for Knowledge Graph Question Answering (KGQA). CogGRAG models the reasoning process as a tree-structured mind map that decomposes the original problem into interrelated subproblems and explicitly encodes their semantic relationships. This structure not only provides a global view to guide subsequent retrieval and reasoning but also enables self-consistent verification across reasoning paths. The framework operates in three stages: (1) top-down problem decomposition via mind map construction, (2) structured retrieval of both local and global knowledge from external Knowledge Graphs (KGs), and (3) bottom-up reasoning with dual-process self-verification. Unlike previous tree-based decomposition methods such as MindMap or Graph-CoT, CogGRAG unifies problem decomposition, knowledge retrieval, and reasoning under a single graph-structured cognitive framework, allowing early integration of relational knowledge and adaptive verification. Extensive experiments demonstrate that CogGRAG achieves superior accuracy and reliability compared to existing methods.
Codehttps://github.com/cy623/RAG.git
Introduction
@@
As a foundational technology for artificial general intelligence (AGI), large language models (LLMs) have achieved remarkable success in practical applications, demonstrating transformative potential across a wide range of domains (llama2; llama3; qwen2). Their ability to process, generate, and reason with natural language has enabled significant advancements in areas such as machine translation (zhu2023multilingual), text summarization (basyal2023text), and question answering (pan2306unifying). Despite their impressive performance, LLMs still face significant limitations in knowledge integration beyond their pre-trained data boundaries. These limitations often lead to the generation of plausible but factually incorrect responses, a phenomenon commonly referred to as hallucinations, which undermines the reliability of LLMs in critical applications.
@@
To mitigate hallucinations, Retrieval Augmented Generation (RAG) (niu2023ragtruth; gao2023retrieval; graphrag; xin2024atomr; chu2024beamaggr; cao2023probabilistic; wu2025pa; zhao20252graphrag) has emerged as a promising paradigm, significantly improving the accuracy and reliability of LLM-generated contents through the integration of external knowledge. However, while RAG successfully mitigates certain aspects of hallucination, it still exhibits inherent limitations in processing complex relational information. As shown in Figure 1 (a), the core limitation of standard RAG systems lies in their reliance on vector-based similarity matching, which processes knowledge segments as isolated units without capturing their contextual interdependencies or semantic relationships (graph_cot). Consequently, traditional RAG implementations are inadequate for supporting advanced reasoning capabilities in LLMs, particularly in scenarios requiring complex problem-solving, multi-step inference, or sophisticated knowledge integration.
@@
Recently, graph-based RAG (graphrag; tog2; graph_cot; gnnrag; xiong2024interactive) has been proposed to address the limitations of conventional RAG systems by incorporating deep structural information from external knowledge sources. These approaches typically utilize KGs to model complex relation patterns within external knowledge bases, employing structured triple representations (<entity, relation, entity>) to integrate fragmented information across multiple document segments. While graph-based RAG has shown promising results in mitigating hallucination and improving factual accuracy, several challenges remain unresolved:
@@
$\bullet$ Lack of holistic reasoning structures. Complex problems cannot be resolved through simple queries; they typically require multi-step reasoning to derive the final answer. Existing methods often adopt iterative or sequential inference pipelines that are prone to cascading errors due to the absence of global reasoning plans (graph_cot; gnnrag). As shown in Figure 1 (b), each step in the iterative process relies on the result of the previous step, indicating that errors occurred at previous steps can propagate to subsequent steps.
@@
$\bullet$ Absence of verification mechanisms. Despite the integration of external knowledge sources, LLMs remain prone to generating inaccurate or fabricated responses when confronted with retrieval errors or insufficient knowledge coverage. Most approaches do not incorporate explicit self-verification, making them vulnerable to retrieval errors or spurious reasoning paths.
To address these challenges, we propose CogGRAG, a Cognition inspired Graph RAG framework, designed to enhance the complex problem-solving capabilities of LLMs in Knowledge Graph Question Answering (KGQA) tasks. CogGRAG is motivated by dual-process theories in human cognition, where problem solving involves both structured decomposition and iterative reflection. Our key contributions are as follows:
$\bullet$ Human-inspired decomposition. CogGRAG introduces a top-down question decomposition strategy that constructs a tree-structured mind map, explicitly modeling semantic relationships among subproblems. This representation enables the system to capture dependencies, causal links, and reasoning order between subcomponents, providing a global structure that guides subsequent retrieval and reasoning.
$\bullet$ Hierarchical structured retrieval. CogGRAG performs both local-level (entity/triple-based) and global-level (subgraph) retrieval based on the mind map. This dual-level retrieval strategy ensures precise and context-rich knowledge grounding.
$\bullet$ Self-verifying reasoning. Inspired by cognitive self-reflection, CogGRAG employs a dual-LLM verification mechanism. The reasoning LLM generates answers, which are then reviewed by a separate verifier LLM to detect and revise erroneous outputs. This design reduces hallucinations and improves output faithfulness.
We validate CogGRAG across three general KGQA benchmarks (HotpotQA, CWQ, WebQSP) and one domain-specific KGQA dataset (GRBench). Empirical results demonstrate that CogGRAG consistently outperforms strong baselines in both accuracy and hallucination suppression, under multiple LLM backbones.
Related Work
@@
Figure 2: The overall process of CogGRAG. Given a target question $Q$, CogGRAG first prompts the LLM to decompose it into a hierarchy of sub-problems in a top-down manner, constructing a structured mind map. Subsequently, CogGRAG prompts the LLM to extract both local level (entities and triples) and global level (subgraphs) key information from these questions. Finally, CogGRAG guides the LLM to perform bottom-up reasoning and verification based on the mind map and the retrieved knowledge, until the final answer is derived.
@@
Reasoning with LLM Prompting. Recent advancements in prompt engineering have demonstrated that state-of-the-art prompting techniques can significantly enhance the reasoning capabilities of LLMs on complex problems (prompt_cot; prompt_tree; prompt_graph). Chain of Thought (CoT (prompt_cot) explores how generating a chain of thought—a series of intermediate reasoning steps—significantly improves the ability of large language models to perform complex reasoning. Tree of Thoughts (ToT) (prompt_tree) introduces a new framework for language model inference that generalizes the popular Chain of Thought approach to prompting language models, enabling exploration of coherent units of text (thoughts) as intermediate steps toward problem-solving. The Graph of Thoughts (GoT) (prompt_graph) models the information generated by a LLM as an arbitrary graph, enabling LLM reasoning to more closely resemble human thinking or brain mechanisms. However, these methods remain constrained by the limitations of the model’s pre-trained knowledge base and are unable to address hallucination issues stemming from the lack of access to external, up-to-date information.
@@
Knowledge Graph Augmented LLM. KGs (wikidata) offer distinct advantages in enhancing LLMs with structured external knowledge. Early graph-based RAG methods (graphrag; G-retriever; GRAG; wu2023retrieve; jin2024survey) demonstrated this potential by retrieving and integrating structured, relevant information from external sources, enabling LLMs to achieve superior performance on knowledge-intensive tasks. However, these methods exhibit notable limitations when applied to complex problems. Recent advancements have introduced methods like chain-of-thought prompting to enhance LLMs’ reasoning on complex problems (Think_on_graph; tog2; graph_cot; chen2024plan; luo2025gfm; luo2024graph; li2024simple). Think-on-Graph (Think_on_graph) introduces a new approach that enables the LLM to iteratively execute beam search on the KGs, discover the most promising reasoning paths, and return the most likely reasoning results. ToG-2 (tog2) achieves deep and faithful reasoning in LLMs through an iterative knowledge retrieval process that involves collaboration between contexts and the KGs. GRAPH-COT (graph_cot) propose a simple and effective framework to augment LLMs with graphs by encouraging LLMs to reason on the graph iteratively. However, iterative reasoning processes are prone to error propagation, as errors cannot be corrected once introduced. This leads to error accumulation and makes it difficult to reason holistically or revise previous steps. In contrast, our CogGRAG framework constructs a complete reasoning plan in advance through top-down decomposition, forming a mind map that globally guides retrieval and reasoning. None of the aforementioned approaches incorporate a self-verification phase. Once an incorrect inference is made, it propagates without correction. CogGRAG addresses this by introducing a dual-LLM self-verification module inspired by dual-process cognitive theory, allowing the system to detect and revise incorrect outputs during the reasoning phase.
Methodology
In this section, we introduce CogGRAG  a human cognition-inspired graph-based retrieval-augmented generation framework designed to enhance complex reasoning in KGQA tasks. CogGRAG simulates human cognitive strategies by decomposing complex questions into structured sub-problems, retrieving relevant knowledge in a hierarchical manner, and verifying reasoning results through dual-process reflection. The overall framework consists of three stages: (1) Decomposition, (2) Structured Retrieval, and (3) Reasoning with Self-Verification. The workflow is illustrated in Figure 2.
Problem Formulation
@@
Given a natural language question $Q$, the goal is to generate a correct answer $A$ using an LLM $p_{\theta}$ augmented with a KG $\mathcal{G}$. CogGRAG aim to design a graph-based RAG framework with an LLM backbone $p_{\theta}$ to enhance the complex problem-solving capabilities and generate the answer $A$.
Top-Down Decomposition
Inspired by human cognition in problem-solving, we simulate a top-down analytical process by decomposing the original question into a hierarchy of semantically coherent sub-questions. This decomposition process produces a reasoning mind map, which serves as an explicit structure to guide subsequent retrieval and reasoning steps. In this mind map, each node represents a sub-question, and edges capture the logical dependencies among them. Specifically, CogGRAG decomposes the original question $Q$ into a hierarchical structure forming a mind map $\mathcal{M}$. Each node in $\mathcal{M}$ is defined as a tuple $m=(q,t,s)$, where $q$ denotes the sub-question, $t$ indicates its depth level in the tree, and $s\in\{\texttt{Continue},\texttt{End}\}$ specifies whether further decomposition is required.
The decomposition proceeds recursively using the LLM:
$\{(q^{t+1}_{j},s^{t+1}_{j})\}_{j=1}^{N}=\texttt{Decompose}(q^{t},p_{\theta},\texttt{prompt}_{dec}),$ (1)
where $N$ is determined adaptively by LLM. This process continues until all leaves in $\mathcal{M}$ are labeled with End, representing atomic questions.
This hierarchical planning mechanism allows CogGRAG to surface latent semantic dependencies that may be overlooked by traditional matching-based RAG methods. For instance, consider the query: “The football manager who recruited David Beckham managed Manchester United during what timeframe?” A conventional retrieval pipeline may fail to resolve the intermediate entity “Alex Ferguson,” as it is not directly mentioned in the input question. In contrast, CogGRAG first decomposes the query into “Who recruited David Beckham?” and “When did that manager coach Manchester United?”, thereby enabling the system to recover critical missing entities and construct a more accurate reasoning path.
By performing top-down question decomposition before any retrieval or reasoning, CogGRAG explicitly separates planning from execution. This design enables holistic reasoning over the entire problem space, mitigates error propagation, and sets the foundation for effective downstream processing.
Structured Knowledge Retrieval
Once the mind map $\mathcal{M}$ has been constructed through top-down decomposition, CogGRAG enters retrieval stage. This phase is responsible for identifying the external knowledge required to support reasoning across all sub-questions. In contrast to prior iterative methods (graph_cot) that retrieve relevant content at each reasoning step, CogGRAG performs a one-pass, globally informed retrieval guided by the full structure of $\mathcal{M}$. This design not only enhances retrieval completeness and contextual coherence, but also avoids error accumulation caused by sequential retrieval failures.
To retrieve relevant information from the KG $\mathcal{G}$, CogGRAG extracts two levels of key information from $\mathcal{M}$. Local-level information captures entities and fine-grained relational facts associated with individual sub-questions. These include entity mentions, entity-relation pairs, and triples. Global-level information reflects semantic dependencies across multiple sub-questions and is expressed as interconnected subgraphs. These subgraphs encode logical chains or joint constraints necessary to resolve the overall problem. For example, given the decomposed question “Which football manager recruited David Beckham?”, we can extract local-level information such as the entity (David Beckham) and the triple (manager, recruited, David Beckham). By additionally considering another decomposed question—such as “During which years did that manager lead Manchester United?”—we can derive global level information, which is represented in the form of a subgraph [(manager, recruited, David Beckham), (manager, manage, Manchester United)]. The extraction process is performed using the LLM:
$\mathcal{K}=\texttt{Extract}(\mathcal{M},p_{\theta},\texttt{prompt}_{ext}),$ (2)
where $\mathcal{K}$ denotes the set of all extracted keys, including entity nodes, triples, and subgraph segments. This design ensures that both isolated facts and relational contexts are captured prior to KGs querying. These keys guide graph-based retrieval. For each entity $e$ in the extracted key set $\mathcal{K}$, we expand to its neighborhood in $\mathcal{G}$, yielding a candidate triple set $\tilde{\mathcal{T}}$. A semantic similarity filter is applied to retain only relevant triples:
$\mathcal{T}=\{\tau\in\tilde{\mathcal{T}},k\in\mathcal{K}\mid\texttt{sim}(\tau,k)>\varepsilon\},$ (3)
where $\texttt{sim}(\cdot)$ is cosine similarity, $\varepsilon$ is a threshold, $\tau$ is the triple in candidate triple set $\tilde{\mathcal{T}}$ and $k$ is the triple in the extracted key set $\mathcal{K}$. The final retrieved triple set $\mathcal{T}$ serves as the input evidence pool for the reasoning module. By leveraging both local and global context from $\mathcal{M}$, CogGRAG ensures that $\mathcal{T}$ is semantically rich, structurally connected, and targeted toward the entire reasoning path.
Reasoning with Self-Verification
The final stage of CogGRAG emulates human problem-solving behavior, where conclusions are not only derived through reasoning but also reviewed through self-reflection (frederick2005cognitive). To this end, CogGRAG employs a dual-process reasoning architecture, composed of two distinct LLMs:
  • LLMres: responsible for answer generation via bottom-up reasoning over the mind map $\mathcal{M}$;
  • LLMver: responsible for evaluating the validity of generated answers based on context and prior reasoning history.
This dual-agent setup is inspired by dual-process theories in cognitive psychology (vaisey2009motivation), where System 1 performs intuitive reasoning and System 2 monitors and corrects errors.
Bottom-Up Reasoning.
Given the final retrieved triple set $\mathcal{T}$ and the structured mind map $\mathcal{M}$, CogGRAG performs reasoning in a bottom-up fashion: sub-questions at the lowest level are answered first, and their verified results are recursively used to resolve higher-level nodes.